Please help! Thermodynamics
Required information NOTE: This is a multi-part question, Once an answer is submitted, you will be unable to return to this part A refrigeration system is to cool bread loaves with an average mass of 430 g from 30°C to -10°C at a rate of 1200 loaves per hour with refrigerated air at -30C. Take the average specific and latent heats of bread to be 2.93 kJ/kg C and 1093 kJ/kg. Determine the rate of heat removal from the breads, in kJ/h. (You must provide an answer before moving to the next part.) The rate of removal of heat from the breads is 116874O kJ/h
media%2Fa21%2Fa2199e84-a8a5-41c2-a0f3-b0
Answer #1

I undrstand that i need to solve only 2nd part as you have already got an answer for the first part.

Heat released = mC*Delta T

m = massof air, C = heat capaity of air = 1 kJ/Kg.K (from literature)

Delta T = 8

116874 = m*1*8 or m = 14609.25

now air density = 1.4282 at average tmeprature of -26 (average of -30 and -22)

volume = mass/density = 14609.25/1.4282 = 10229.13 m3/h